Could you show me how to sove this using the elimination method?
7x+5y=2
8x-9y=17

Answer by checkley71(8403)

ELIMINATE THE Y TERM BY MULTIPLYING EQ #1 BY 9 & EQ #2 BY 5 & THEN ADD THEM
63X+45Y=18
40X-45Y=85
-------------
103X=103 OR X=103/103 OR X=1 SUBSTITUTING 1FOR X IN EITHER EQUATION TO FIND Y
7*1+5Y=2 OR 7+5Y=2 OR 5Y=-5 OR Y=-5/5 OR Y=-1
PROOF 8*1-9(-1)=17 OR 8+9=17 OR 17=17